OpenStudy (anonymous):

Which two options could be the next step in simplifying this expression? (18 + 23) + 7 A. 41 + 7 (use order of operations) B. 30 + 25 (use the commutative property) C. 25 + 30 (use the distributive property) D. 18 + 30 (use the associative property)

OpenStudy (inkyvoyd):

hint, order of operations only applies when there's more than one type of operation, commutative means a operator b = b operator a, distributive needs 2 operators, and associative needs at least 3 numbers with the same operation

OpenStudy (jonathanvaldes98):

i think A.

OpenStudy (calculusxy):

Do you know what all of the math vocabulary means in the parentheses?

OpenStudy (anonymous):

The answer is A and D. (18 + 23) + 7 = 41 + 7 18 + (23 + 7) = 18 + 30

OpenStudy (jonathanvaldes98):



OpenStudy (ybarrap):

Do parenthesis 1st then the rest, A would be best answer.

OpenStudy (calculusxy):

But aren't they asking for two options meaning two answer choices?

OpenStudy (anonymous):

Thanks!

OpenStudy (debbieg):

@s01025881, the problem is it says the NEXT step. As you point out, D could be arrived at, but it takes the associative property + 1 additional step: 18 + (23 + 7) Associative = 18 + 30 Order of operations The 18 + 30 is not the result of the associative property, but the step after that.

OpenStudy (anonymous):

so what is the answer?